OpenStudy (anonymous):

How to solve for x with steps? 3x+7=x

OpenStudy (anonymous):

3x+7=x 3x+7-x=0 2x+7=0 x=-7/2

OpenStudy (anonymous):

Can you explain further the answer is kinda messy.

OpenStudy (anonymous):

ok see here : first of all transpose x to left hand side \[3x-x+7=0\] \[ 2x+7=0\] tranpose 7 to right hand side \[2x=0-7\] \[2x=-7\] divide both sides by 2 \[2x/2 = -7/2\] \[x=-7/2\]

OpenStudy (anonymous):

i hope u must have got it now. any other problem u can ask it now

OpenStudy (anonymous):

x/5 - 12 =7

OpenStudy (anonymous):

bmanny , which grade are you in

OpenStudy (anonymous):

ok ask it in the left box now

OpenStudy (anonymous):

my answer was also correct, i should also be given medal ;-)

OpenStudy (anonymous):

bmanny5 r u here

OpenStudy (anonymous):

thanks bmanny ..

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

bmanny , which grade are you studying

OpenStudy (anonymous):

ask it there in the left box now

OpenStudy (anonymous):

have u asked it

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

bmanny ap kaun si class mein padte ho

OpenStudy (anonymous):

csr21 u r indian